OpenStudy (anonymous):

Simplify x^1/3 * x^1/4 x^1/7 x^1/12 x^7/12 2x^7/12

OpenStudy (anonymous):

when you multply two terms, add the exponents... but u need common denominators..(12)

OpenStudy (anonymous):

4/12+3/12=7/12

OpenStudy (anonymous):

hint.. \[\LARGE \sqrt[3]{x}=(x)^{\frac13}\] and \[\LARGE x^a\cdot x^b=x^{a+b}\] use them...